• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Zolnierczuk

    Log entry time 08:15:39 on August10,2001

    Entry number 70280

    keyword=owl shift summary

    0:00 configuration: target 270deg (l/2 out), beam l/2 out

    0:00 Run 2741/21741 2.4/2.3M
    0:45 Run 2742/21742 2.9/2.8M
    1:26 Run 2743/21743 2.7/2.6M
    2:06 Run 2744/21744 2.9/2.8M
    2:47 Run 2745/21745
    3:30 CODA problems, TS0/Right arm rebooted
    R-arm datamon froze, tried to restart it, but failed
    3:45 Run 2747/21747 ?/?M
    4:20 still CODA problems
    4:40 NMR/EPR done (35.2/36.7)
    5:00 called Bob - it turned out to be scalServ problem
    on TS0/R-arm, after few resets and some cursing
    we're back in business
    5:40 Run 2757/21757 2.8/2.7M
    6:19 Run 2758/21758 2.0/1.9M
    7:00 Run 2759/21759 2.6/2.5M
    7:43 Run 2760/21760 still running